At certain place, the horizontal component of earths magnetic field is 3G and the angle of dip at the place is 30 degree. The magnetic field of earth at that location is
1
Expert's answer
2019-06-03T09:55:03-0400


We can find the magnetic field of the Earth from the geometry:


"cos \delta = \dfrac{B_h}{B},"

here, "\delta = 30^{\circ}" is the angle of dip, "B_h = 3G" is the horizontal component of Earth's magnetic field, "B" is the magnetic field of the Earth.

Then, from this formula we can find the magnetic field of the Earth:


"B = \dfrac{B_h}{cos \delta} = \dfrac{3G}{cos 30^{\circ}} = 3.46G."

Answer:

"B = 3.46G."